A system is described by the following transfer function:
$2 + 16 H(s) = - 11s(s+2)(2+1)

A) What is the frequency response, H(f)?

B) What is the magnitude and phase (in degrees) of the frequency response at a frequency of w=3 rads/sec, corresponding to f= \frac{3}{2\pi } hz?
Answer #1

316 - رو) slet (stis ! (+ ۔ ارى لمر = | مارک/ و أ+ (54 ) ۔۔ الم ( 114= ه (217f;) (LITf+z) T- CLIFF رم - ها, ے اماه - 16 (أ+ 2